# IEEEexample:BSTcontrol

### v1

@@ Abstract
We study parameterizations of stabilizing nonlinear policies for learning-based control. We propose a structure based on a nonlinear version of the Youla-Kučera parameterization combined with robust neural networks such as the recurrent equilibrium network (REN). The resulting parameterizations are unconstrained, and hence can be searched over with first-order optimization methods, while always ensuring closed-loop stability by construction. We study the combination of (a) nonlinear dynamics, (b) partial observation, and (c) incremental closed-loop stability requirements (contraction and Lipschitzness). We find that with any two of these three difficulties, a contracting and Lipschitz Youla parameter always leads to contracting and Lipschitz closed loops. However, if all three hold, then incremental stability can be lost with exogenous disturbances. Instead, a weaker condition is maintained, which we call d-tube contraction and Lipschitzness. We further obtain converse results showing that the proposed parameterization covers all contracting and Lipschitz closed loops for certain classes of nonlinear systems. Numerical experiments illustrate the utility of our parameterization when learning controllers with built-in stability certificates for: i) “economic” rewards without stabilizing effects; ii) short training horizons; and iii) uncertain systems.
@@ 1.1 Feedback Policy Parameterizations
One way to take advantage of prior knowledge is to design a “base controller” — for example, using classical model-based techniques — and augment it with a neural network as in Fig. 1(b). This is called residual RL [12, 13, 14]. The base controller is used to bootstrap performance and allows us to naturally fuse model-based control with data-driven RL. However, closed-loop stability and robustness constraints are typically not imposed in residual RL. It would be possible to do so via (for example) the small gain theorem if a gain bound from the neural network output to the system output is known and gain-bounded neural networks are employed. However, this will typically be very conservative.
@@ 3 Nonlinear Youla Parameterization
Figure 2: A version of the Youla-Kucera parameterization. The Youla parameter $\mathcal{Q}$ augments an observer-based feedback controller $\mathcal{K}_{b}$ to control a plant $\mathcal{G}$. The known exogenous inputs $\eta$ have been omitted to simplify the diagram.
@@ 4.1 Restricting to LTI Systems
The Youla controller (12) therefore parameterizes all contracting and Lipschitz closed loops for a given LTI system.
@@ 4.2 Restricting to Perfect State Feedback
As with the LTI setting, the Youla controller (18) with innovations (19) parameterizes all contracting and Lipschitz closed-loops of (17). This formulation has strong connections to disturbance-feedback parameterizations used in MPC [21]. Similar results to Theorem 2 in terms of finite $\ell_{p}$-gain stability have also been derived in the context of systems-level synthesis (SLS) [51, 46].
@@ 4.3 Restricting to Finite-Gain Stability
Moving to the partially-observed nonlinear setting, similar results have been proven for the weaker notion of finite-gain stability (e.g., [29, 30]). We summarize these results for our specific problem setup below. Let us first make the following assumptions.
@@
Comparing to the partially-observed LTI setting, assumptions A1A3 are similar to assuming that we can design a stabilizing linear base controller (14). Importantly, A2 and A3 imply that the observer error will always exponentially decay in the absence of disturbances $d$, regardless of the control inputs $u$ [11, Prop. 2].
Under the above assumptions, the Youla parameterization achieves the following result (similar to e.g., [29, 30]).
@@
    Any controller $u=\mathcal{K}(\eta,y)$ of the form (16) achieving a finite-gain stable closed loop can be written in the form (12) with finite-gain stable $\mathcal{Q}$.
This implies that the Youla parameterization represents all finite-gain stabilizing controllers for a given partially-observed nonlinear system.
5 Results for Partially-Observed Nonlinear Systems & Incremental Stability
@@ 6 Converse Results
In Section 4 we saw that, in certain simplified settings, the proposed Youla parameterization includes all and only controllers resulting in contracting and Lipschitz closed-loops. In Section 5 we saw by counterexample that this is not true in the partially-observed nonlinear case, but that the weaker property of d-tube contraction and Lipschitzness holds. This raises the question: what class of controllers are covered by the Youla parameterization? We provide an answer in this section, and include proofs in Appendix .6.
@@ 7.3 Long-Term Stability with Short-Term Training
with $q>0$. The optimal LQG controller has arbitrarily small gain and phase margins as $\sigma_{w},q$ increase [56]. This makes the problem challenging in a deep RL setting, since optimal and unstable policies may appear very close in parameter space.
@@ 8 Conclusions
We have studied a nonlinear version of the Youla parameterization for nonlinear dynamical systems, and presented a constructive neural realization called the Youla-REN. We have proven that: the parameterization guarantees contracting and Lipschitz closed-loops when the innovations signal is decoupled from the Youla augmentation; if these signals are coupled due to additive disturbances, the disturbed trajectories are constrained within bounded tubes centered on the disturbed-free trajectories; and we cover all contracting and Lipschitz closed loops for certain nonlinear systems. Using RENs for the Youla parameter, empirical studies show that we can learn stabilizing and performant Youla-REN policies on feedback control tasks, even if they are trained with arbitrary (potentially de-stabilizing) cost functions, over short time horizons, and with model uncertainty. For future work, we aim to leverage these advantages and apply Youla-based policy parameterizations to real-world robotic control.
@@
We now show that the effect of disturbances in constrained within a tube of bounded $\ell_{2}$ norm about the disturbance-free trajectories. Let $\xi_{t}=[x_{t};s_{t}]\in\mathbb{R}^{n_{x}+n_{s}}$ and $q_{t}\in\mathbb{R}^{n_{q}}$ be the states of the base-controlled closed-loop system $(\eta,d,\tilde{u})\mapsto z$ and the Youla parameter $(\eta,\tilde{y})\mapsto\tilde{u}$, respectively. For convenience we denote the state of the total closed-loop system by $\bar{x}_{t}=[\xi_{t};\hat{x}_{t},q_{t}]$. Consider two copies of the system starting at initial states $\bar{x}_{0},\bar{x}_{0}^{\star}$ with inputs $(\eta,w,v)\in\ell$ and $(\eta^{\star},0,0)\in\ell$, respectively, where $x^{\star}_{0}=\hat{x}_{0}^{\star}$. From Assumption (A1) and Lipschitzness of the Youla parameter respectively, we have

### v2

@@ Abstract
We study parameterizations of stabilizing nonlinear policies for learning-based control. We propose a structure based on a nonlinear version of the Youla-Kučera parameterization combined with robust neural networks such as the recurrent equilibrium network (REN). The resulting parameterizations are unconstrained, and hence can be searched over with first-order optimization methods, while always ensuring closed-loop stability by construction. We study the combination of (a) nonlinear dynamics, (b) partial observation, and (c) incremental closed-loop stability requirements (contraction and Lipschitzness). We find that with any two of these three difficulties, a contracting and Lipschitz Youla parameter always leads to contracting and Lipschitz closed loops. However, if all three hold, then incremental stability can be lost with exogenous disturbances. Instead, a weaker condition is maintained, which we call d-tube contraction and Lipschitzness. We further obtain converse results showing that the proposed parameterization covers all contracting and Lipschitz closed loops for certain classes of nonlinear systems. Numerical experiments illustrate the utility of our parameterization when learning controllers with built-in stability certificates for: (i) “economic” rewards without stabilizing effects; (ii) short training horizons; and (iii) uncertain systems.
@@ 1.1 Feedback Policy Parameterizations
One way to take advantage of prior knowledge is to design a “base controller” — for example, using classical model-based techniques — and augment it with a neural network as in Fig. 1(b). This is called residual RL [12, 13, 14]. The base controller is used to bootstrap performance and allows us to naturally fuse model-based control with data-driven RL. However, closed-loop stability and robustness constraints are typically not imposed in residual RL. It would be possible to do so via (for example) the small-gain theorem if a gain bound from the neural network output to the system output is known and gain-bounded neural networks are employed. However, this will typically be very conservative.
@@ 3 Nonlinear Youla Parameterization
Figure 2: A version of the Youla-Kučera parameterization. The Youla parameter $\mathcal{Q}$ augments an observer-based feedback controller $\mathcal{K}_{b}$ to control a plant $\mathcal{G}$. The known exogenous inputs $\eta$ have been omitted to simplify the diagram.
@@ 4.1 Restricting to LTI Systems
The Youla parameterization represents all and only contracting and Lipschitz closed loops for a given LTI system.
@@ 4.2 Restricting to Perfect State Feedback
As with the LTI setting, the Youla parameterization represents all and only contracting and Lipschitz closed-loops with (17). This formulation has strong connections to disturbance-feedback parameterizations in MPC [21]. Similar results to Theorem 2 in terms of finite-gain stability have also been derived for system-level synthesis (SLS) [51, 46].
@@ 4.3 Restricting to Finite-Gain Stability
Moving to the partially-observed nonlinear setting, similar results have been proven for the weaker notion of finite-gain stability (e.g., [32, 29]). We summarize these results for our specific problem setup given the following assumptions.
@@
Comparing to the partially-observed LTI setting, assumptions A1A3 are similar to assuming that we can design a stabilizing linear controller (14). Importantly, A2A3 imply that the observer error exponentially decays when there are no disturbances, regardless of the control inputs [11, Prop. 2].
Under the above assumptions, the Youla parameterization achieves the following result (similar to e.g., [29, Prop. 21], [32, Thm. 3.1]).
@@
    If the base controller (8) is finite-gain stable, then any controller $u=\mathcal{K}(\eta,y)$ of the form (16) achieving a finite-gain stable closed loop can be written in the form (12) with finite-gain stable $\mathcal{Q}$.
This implies that the Youla parameterization represents all and only finite-gain stabilizing controllers for a given partially-observed nonlinear system.
5 Results for Partially-Observed Nonlinear Systems with Incremental Stability
@@ 6 Converse Results
In Section 4 we saw that, in certain simplified settings, the proposed Youla parameterization includes all possible and only those controllers resulting in contracting and Lipschitz closed-loops. In Section 5 we saw by counterexample that this is not true in the partially-observed nonlinear case, but that the weaker property of d-tube contraction and Lipschitzness holds. This raises the question: what class of controllers are covered by the Youla parameterization? We provide an answer in this section, and include proofs in Appendix .6.
@@ 7.3 Long-Term Stability with Short-Term Training
with $q>0$. The optimal LQG controller has arbitrarily small-gain and phase margins as $\sigma_{w},q$ increase [56]. This makes the problem challenging in a deep RL setting, since optimal and unstable policies may appear very close in parameter space.
@@ 8 Conclusions
We have studied a nonlinear version of the Youla parameterization for nonlinear dynamical systems, and presented a constructive neural realization called the Youla-REN. We have proven that: the parameterization guarantees contracting and Lipschitz closed loops when the innovations signal is decoupled from the Youla augmentation; if these signals are coupled due to additive disturbances, the disturbed trajectories are constrained within bounded tubes centered on the disturbed-free trajectories; and we cover all contracting and Lipschitz closed loops for certain nonlinear systems. Using RENs for the Youla parameter, empirical studies show that we can learn stabilizing and performant Youla-REN policies on feedback control tasks, even if they are trained with arbitrary (potentially de-stabilizing) cost functions, over short time horizons, and with model uncertainty. For future work, we aim to leverage these advantages and apply Youla-based policy parameterizations to real-world robotic control.
@@
We now show that the effect of disturbances in constrained within a tube of bounded $\ell_{2}$ norm about the disturbance-free trajectories. Let $\xi_{t}=[x_{t};s_{t}]\in\mathbb{R}^{n_{x}+n_{s}}$ and $q_{t}\in\mathbb{R}^{n_{q}}$ be the states of the base-controlled closed-loop system $(\eta,d,\tilde{u})\mapsto z$ and the Youla parameter $(\eta,\tilde{y})\mapsto\tilde{u}$, respectively. For convenience we denote the state of the total closed-loop system by $\bar{x}_{t}=[\xi_{t};\hat{x}_{t};q_{t}]$. Consider two copies of the system starting at initial states $\bar{x}_{0},\bar{x}_{0}^{\star}$ with inputs $(\eta,w,v)\in\ell$ and $(\eta^{\star},0,0)\in\ell$, respectively, where $x^{\star}_{0}=\hat{x}_{0}^{\star}$. From assumption (A1) and Lipschitzness of the Youla parameter respectively, we have